# MiM: Mask in Mask Self-Supervised Pre-Training for 3D Medical Image Analysis

### v1

@@ Abstract
The Vision Transformer (ViT) has demonstrated remarkable performance in Self-Supervised Learning (SSL) for 3D medical image analysis. Mask AutoEncoder (MAE) for feature pre-training can further unleash the potential of ViT on various medical vision tasks. However, due to large spatial sizes with much higher dimensions of 3D medical images, the lack of hierarchical design for MAE may hinder the performance of downstream tasks. In this paper, we propose a novel Mask in Mask (MiM) pre-training framework for 3D medical images, which aims to advance MAE by learning discriminative representation from hierarchical visual tokens across varying scales. We introduce multiple levels of granularity for masked inputs from the volume, which are then reconstructed simultaneously ranging at both fine and coarse levels. Additionally, a cross-level alignment mechanism is applied to adjacent level volumes to enforce anatomical similarity hierarchically. Furthermore, we adopt a hybrid backbone to enhance the hierarchical representation learning efficiently during the pre-training. MiM was pre-trained on a large scale of available 3D volumetric images, i.e., Computed Tomography (CT) images containing various body parts. Extensive experiments on thirteen public datasets demonstrate the superiority of MiM over other SSL methods in organ/lesion/tumor segmentation and disease classification. We further scale up the MiM to large pre-training datasets with more than 10k volumes, showing that large-scale pre-training can further enhance the performance of downstream tasks. The improvement also concluded that the research community should pay more attention to the scale of the pre-training dataset towards the healthcare foundation model for 3D medical images.
keywords:
[…]
journal: Preprint
@@ 1 Introduction
Fig. 1: Different SSL for 3D medical image analysis. Current Masked Image Modeling methods for 3D medical images primarily (a) rely on pretext tasks e.g., inpainting, at a single level, utilizing hybrid transformers to incorporate all tokens or (b) employ an MAE that reconstructs at a single level using visible tokens. In contrast, (c) we observe that 3D medical images inherently exhibit hierarchical properties. Thus, our Mask in Mask (MiM) framework aims to encode multi-level 3D medical image learning across hierarchical visual tokens at various scales through multi-level reconstruction and cross-level alignment (we set the number of level $L$ to 3 in this figure). Additionally, our framework employs a hybrid transformer while only using visible tokens.
Deep learning has made significant progress in medical image analysis, albeit limited to the supervised learning paradigm that relies on a relatively large amount of labeled training data [57, 93]. However, labeling 3D medical images is a time-consuming and labor-intensive process that requires expertise [39, 60]. Specifically, to expand the training set while reducing human intervention, Semi-supervised Learning and Self-Supervised Learning (SSL) has been become a promising direction  [16, 30, 31, 10, 74]. Among them, SSL has shown great potential in learning transferable representations from unlabeled data without any annotations, which can be further fine-tuned on downstream tasks to enhance the task performance [38, 34, 62, 78, 70, 91, 89, 94].
MAE-based methods [30, 6, 64, 21, 66, 67, 46, 79] has achieved high performance in various natural image analysis tasks by learning transferable representations through reconstructing the original output in pixels from its highly-masked version based on the ViT structure. However, 3D medical images such as Computed Tomography (CT), which are often volumetric and in a much higher dimension, contain a multi-granularity of visual details. Existing methods [38, 34, 62, 78, 70, 91, 89, 94, 76, 69, 72] like Swin-UNETR[62] proposes to learn local and global representation by combining tailored proxy tasks including inpainting, rotation, and contrastive different views. He et al. [34] further explores the geometric similarity between global and local crops by affirmation and matching. Jiang et al. [38], on the other hand, maximizes the similarity between different views that contain consistent semantics. MAE3D [16] reconstructs the missing pixel based on a cropped volume. These methods mainly focus on encoding the 3D medical image from local or global perspectives at a single scale, which may still result in the loss of finer granularity detail.
In the case of medical images in large spatial sizes of high dimensions, the use of hierarchical design is crucial for effective analysis and interpretation of clinical information [9, 62, 34, 35]. In this paper, we propose a novel framework that enhances representation learning of MAE for 3D medical images by hierarchical design, as illustrated in Fig. 1. To achieve this goal, our framework utilizes a hierarchical view of 3D medical images, containing rich multi-granularity of clinical information essential for further analysis and providing important clinical insights. To obtain hierarchical masked volumetric inputs, we propose a multi-level volume generation strategy. This strategy conveys information from a coarse to a finer level, which combines to bring forth a rich representation of the model. We employ multi-level reconstruction to simultaneously reconstruct the masked volumes with multi-granularity details to learn discriminative representation. Compared to single-level reconstruction [16, 66, 79, 22], multi-level reconstruction can capture more detailed information and improve representation learning [67]. Our multi-levels are constructed by dividing the original volume into multiple levels, which is different from the multi-scale strategy that uses different scales of the same volume [67, 35]. Cross-level alignment between adjacent level volumes is applied to learn both the global context information and local details, achieving better representation. Positive pairs i.e., volumes from the same original volumes but at the adjacent level granularity, share the same clinical information. We enable the model to learn more robust representations by aligning the features of these volumes at different scales. We extend the hybrid backbone for MAE [22] from 2D natural images to our 3D medical image pre-training framework to further improve the efficiency during the pre-training period. Since previous MAE-based works [79] require all tokens to be kept in the later transformer stages, leading to an unaffordable computation cost. This leads to the loss of MAE’s efficiency advantage.
In summary, our contribution can be categorized into three aspects:
@@
    We propose a simple yet effective SSL framework, called MiM, which advances MAE with hierarchical design in 3D medical image pre-training, pushing the representability of pre-trained models.
@@
    MiM encodes multi-level visual cues across various scales into the metric, providing a reliable measurement of the hierarchical structure of 3D medical images. A multi-level volume generation strategy is proposed, and two tailored proxy tasks, multi-level reconstruction, and cross-level alignment, are presented, which simultaneously learn powerful local and global representations.
@@
    We pre-train MiM with various sizes of pre-training datasets ranging from 1k to 10k volumes. Extensive experiments demonstrate its effectiveness and achieve state-of-the-art performance across various tasks on thirteen public datasets. We conclude that scaling up the pre-training dataset can further improve the performance of MiM.
@@ 2 Related works
In this section, we first review the previous mainstream self-supervised learning paradigms, especially for existing SSL methods for 3D medical images. Then we introduce the recent advances in masked image modeling and hierarchical design of SSL for 3D medical images for comparison with our method and highlight the differences.
SSL for Medical Image Analysis. Previous studies in SSL can be categorized into two primary paradigms: Contrastive learning and Pretext task [31, 10, 23]. Contrastive learning methods bring positive pairs of images (e.g., different views of the same image) closer and separate negative pairs (e.g., different images) away in the feature space [11, 10, 31, 12, 15]. For example, MoCo v3 utilized a memory bank to maintain consistent representations of negative samples during contrastive learning [31, 13, 15]. These works have also shown effectiveness in medical image analysis, such as dermatology classification and chest X-ray classification [87, 90, 80]. However, previous contrastive learning approaches have primarily focused on the semantic misalignment of different instances, leading to minor improvements in dense prediction, such as segmentation [5, 75, 84, 90]. Pretext task, on the other hand, are designed to explore the inner space structure information of images and have shown promise in dense prediction tasks [7, 26, 63, 77]. For example, Zhou et al. [91] proposed to directly pre-train a CNN model by restoring volumetric medical images from their distorted versions. Zhu et al. [92] utilized the concept of solving a Rubik’s Cube to learn structural features from the volumetric medical data. Additionally, five common pretext tasks have been verified to be effective for volumetric image pre-training [61]. Tang et al. [62] proposed to pre-train a 3D swin transformer [48] with the combination of three pretext tasks. Recently, Masked Image Modeling [30, 79], considered as widely used pretext tasks based, shows its effectiveness on 3D medical analysis [16, 76, 94].
Masked Image Modeling. Masked Image Modeling is a generative SSL technique that involves training on images that have been corrupted by masking [30, 79, 16, 6, 66, 70, 76, 69, 94]. Previous researches on this technique, such as Vincent et al. [65] and Pathak et al. [52], treated masking as a type of noise and used inpainting techniques to predict the missing pixels. Dosovitskiy et al. [18] investigated masked patch prediction by predicting the mean color of images. He et al. [30] adhered to the spirit of raw pixel restoration and demonstrated that masking a high proportion of input images can yield a non-trivial and meaningful self-supervisory task.  Chen et al. [6], Wang et al. [66] proposes to improve masking priority strategies [66] and the reconstruction targets [6]. Zhang et al. [81], Wang et al. [69] introduce the concept of Masked Image Modeling to versetile 2D medical scenario e.g. disease classfication [76] and segmentation [69]. Chen et al. [16] represented a recent advancement in Masked Image Modeling [30, 79] for 3D CT in medical image analysis. Zhuang et al. [94] proposed to advance MAE for 3D medical image analysis, from the global and local perspectives by incorporating the contrastive learning paradigm.
Our method follows the primary idea of Masked Image modeling. As shown in Fig. 1, to address the existing problems mentioned above, instead of single-level masking, MiM leverages the hierarchical design inherent in the 3D medical images to generate multi-level volume for reconstruction and cross-level alignment, which can effectively capture the hierarchical structure of 3D medical images and improve the representation learning.
Hierarchical SSL. Hierarchical SSL has been investigated in numerous studies [79, 6, 22, 83] in the context of natural images.  Xie et al. [79] and  Chen et al. [6] proposed to incorporat Masked Image Modeing [30] with a hierarchical backbone e.g., Swin Transformer [48]. Wang et al. [67] reconstructing masked patches on multi-scale features using ViT [19], while  Gao et al. [22] advocated for predicting masked patches using only visible patches and a hierarchical backbone. Hosseinzadeh Taher et al. [35] divided single-scale images into multi granularities and aligned their contrastive views [31, 12] with ResNet [32].
However, the relatively low dimensionality of 2D images limits the effectiveness of learning the hierarchical structure of 3D medical images given a single-level image with hybrid features. Furthermore, previous methods [35, 67] treated each level image separately, neglecting the cross-level shared semantics.
In this paper, we propose a novel pretext task of Masked Image Modeling with a hierarchical design tailored for 3D medical images. This approach considers the inherent hierarchical design of 3D medical images during pre-training. Distinct from previous methods, we introduced a totally different Masked Image Modeling paradigm for 3D medical images. Our proposed hierarchical Masked Image Modeling paradigm employs multi-level volume generation for reconstruction and cross-level alignment, effectively capturing the hierarchical structure of 3D medical images and enhancing representation learning.
@@ 3 Methodology
This section provides an overview of our proposed MiM method. Firstly, the overall framework of the MiM method is introduced in Section 3.1. Secondly, the process of multi-level reconstruction is presented in Section 3. Then, the cross-level alignment via contrastive learning in our proposed MiM method is described in Section 3.3. Finally, the backbone of the MiM method in the pre-training period is introduced in Section 3.4.
Fig. 2: The overall framework of our MiM pre-training framework. The level $L$ is set to 3 for better illustration. We first conduct the process of multi-level masked volume generation. The multi-level reconstruction module is responsible for reconstructing the masked volumes at different levels. The cross-level alignment module aligns representations of volumes between volumes from adjacent levels, aiming to enforce anatomical similarity in a hierarchical manner.
@@ 3.1 Overall framework
The proposed MiM method’s overall framework is presented in Fig. 2, which comprises multi-level reconstruction modules and cross-level alignment modules. To pre-train the model with MiM, we first generate multi-level volumes from the input 3D medical images. Then, an input volume is cropped into non-overlapping patches, which are divided into unmasked patches and masked patches. The unmasked patches are transformed into a high-dimension feature space using a typical backbone (CNN [32] and transformer [18]). The masked patches are used to generate the next level of masked volumes for multi-level learning. The goal is to restore the masked patches from different levels of masked volumes. In this paper, instead of a single level of masked volume [21, 66, 67], we propose to ease this goal by multi-level of masked volume. We developed a $\mathcal{L}_{R}$ to supervise the final prediction. In addition, we further use a loss function $\mathcal{L}_{C}$ to align the shared semantics between the cross-level volumes, aiming to learn the global semantic content as well as local details. Further details are presented in Section 3 and Section 3.3.
@@ 3.2 Multi-level reconstruction
Generation of multi-level masked volumes. Given a volumetric image $x\in\mathbb{R}^{CHWD}$ (e.g., $C=1$ for CT), we aim to generate multi-level volumes $\{x^{l}\in\mathbb{R}^{CH^{l}W^{l}D^{l}},l\in L\}$, which contains $L$ numbers of level, i.e., different granularity information from coarse to fine level. The generation of multi-level volume is illustrated in Fig. 3. We start by cropping a large portion of sub-volume from $x$ as Level-1 volume $x^{1}\in\mathbb{R}^{CH^{1}\times W^{1}\times D^{1}}$. The level 1 volume $x^{1}$ is then patchified into $N$ non-overlapping visual tokens using a patch of $\{\frac{H^{1}}{6},\frac{W^{1}}{6},\frac{D^{1}}{6}\}$ as in  [30, 16, 94]. Following previous MAE-based methods [16, 94], we apply a high mask ratio $\mu$ (e.g., 60%) to the $N$ non-overlapping visual tokens (e.g., 216), resulting in visible tokens and invisible tokens. Each visible token is resized at spatial dimension to $\mathbb{R}^{Chwd}$ before further processed by a linear project as previous MAE-based methods [16, 94]. The invisible tokens from Level-1 $x^{1}$ are considered as the next level volume, i.e., Level-2 volume $x^{2}\in\mathbb{R}^{CH^{2}\times W^{2}\times D^{2}}$. It’s noted that the Level-2 volume $x^{2}$ is generated from the masked patches of Level-1 volume $x^{1}$, instead of unmasked patches. The content of the target reconstruction at different levels shares regions, meaning that the volumes for Level-1 and Level-2 reconstruct overlapping regions but from different granularities, respectively. This can effectively capture the hierarchical structure of 3D medical images and improve representation learning [67]. Ablation study of the reconstruction targets, i.e., the choices of masked patches or unmasked patches as next level volume, is presented in Section 4.7.1. Since there exist many invisible tokens, we randomly sample $\gamma$ times without replacement from these tokens for computation efficiency.
@@
Fig. 3: Illustration of the multi-level masked volume generation. The level $L$ is set to 2 and slices drawn from the 3D medical images are shown in the figure for better illustration. Level-2 volume $x^{2}$ is randomly sampled from the masked patches of Level-1 volume $x^{1}$ (i.e., patch with red box). Consequently, the content of the target reconstruction at different levels shares regions, meaning that the volumes for Level-1 and Level-2 reconstruct overlapping regions but from different granularities, respectively.
The Level-2 volume $x^{2}$ is also patchified into $N$ non-overlapping visual tokens with a patch size of $\{\frac{H_{2}}{6},\frac{W_{2}}{6},\frac{D_{2}}{6}\}$. With applying the mask ratio $\mu$ as in the previous process, the visible tokens from Level-2 volume $x^{2}$ are resized to $\mathbb{R}^{Chwd}$ before feeding the backbone. Meanwhile, the invisible tokens from Level-2 volume $x^{2}$ are considered as next finer level volume i.e., Level-3 volume $x^{3}\in\mathbb{R}^{CH^{3}\times W^{3}\times D^{3}}$. Similarly, we randomly sample $\beta$ times without replacement from these tokens for computation efficiency. This generation process of multi-level volume is repeated until the last level of volume $x^{L}$ is generated.
@@
Thus, all visible tokens from different levels are fed into the backbone to extract the high-dimension features $z$. Following previous SSL methods [30, 16, 94], a lightweight decoder is used to project $z$ along with learnable invisible tokens into latent features $q$. For reconstruction, the decoder followed by a simple prediction head is used to reconstruct the invisible tokens $y$. For cross-level alignment, all visual tokens from identical levels are aggregated by MLP layers and then projected into a context vector $c$.
@@
Single-level reconstruction.  Our reconstruction target is the pixel values of the invisible tokens for each level volume. With feature extracted from the backbone and the decoder with a predition head (i.e., a linear layer), following previous MAE methods [16, 94], we reshape the prediction results and reconstruction targets to one-dimensional vector, i.e., $\hat{y}\in\mathbb{R}^{{1\times C}}$ and $y\in\mathbb{R}^{{1\times C}}$, where $C$ is the number of dimensions. Specifically, we empirically set the $C$ to 768, as in  [30, 16, 94].
@@
Then we compute the differences $d$ between the reconstruction target $y$ and the prediction result $\hat{y}$. Specifically, we use MSE distance [30, 16] to measure the difference $d_{m}$ between the prediction result and the reconstruction target for each masked token $m$ as follows:
@@
where $M$ denotes the numbers of masked tokens and $d_{m}$ ranges from 0 to 1. Lower $d_{m}$ indicates the smaller difference between the prediction result $\hat{y}_{m}$ and the reconstruction target $y_{m}$.
Aiming to minimize the difference $d_{m}$, we define the reconstruction loss $\mathcal{L}_{\mathcal{R}}^{l}$ for each level of masked volumetric images $l\in L$ as follows:
@@
where $M$ represents the number of mask tokens. The reconstruction loss is computed as the sum of squared differences between the reconstruction target and the reconstructed representations based on pixel values.
Loss function for multi-level reconstruction. To learn multi-granularity details from 3D volumetric images, We apply multi-level reconstruction for each level of masked volumetric images, ranging from coarser levels to finer levels. The formulation of the multi-level reconstruction loss $\mathcal{L}_{\mathcal{R}}$ is shown as follows:

### v2

@@ Abstract
The Vision Transformer (ViT) has demonstrated remarkable performance in Self-Supervised Learning (SSL) for 3D medical image analysis. Masked AutoEncoder (MAE) for feature pre-training can further unleash the potential of ViT on various medical vision tasks. However, due to large spatial sizes with much higher dimensions of 3D medical images, the lack of hierarchical design for MAE may hinder the performance of downstream tasks. In this paper, we propose a novel Mask in Mask (MiM) pre-training framework for 3D medical images, which aims to advance MAE by learning discriminative representation from hierarchical visual tokens across varying scales. We introduce multiple levels of granularity for masked inputs from the volume, which are then reconstructed simultaneously ranging at both fine and coarse levels. Additionally, a cross-level alignment mechanism is applied to adjacent level volumes to enforce anatomical similarity hierarchically. Furthermore, we adopt a hybrid backbone to enhance the hierarchical representation learning efficiently during the pre-training. MiM was pre-trained on a large scale of available 3D volumetric images, i.e., Computed Tomography (CT) images containing various body parts. Extensive experiments on twelve public datasets demonstrate the superiority of MiM over other SSL methods in organ/tumor segmentation and disease classification. We further scale up the MiM to large pre-training datasets with more than 10k volumes, showing that large-scale pre-training can further enhance the performance of downstream tasks. Code and checkpoint will be available once upon acceptance.
{IEEEkeywords}
CT, Self-Supervised Learning, Segmentation, Classification, 3D medical images.
@@ 1 Introduction
The advent of deep learning has catalyzed unprecedented advances in medical image analysis, predominantly within the supervised learning paradigm. However, this paradigm’s fundamental limitation lies in its reliance on extensively labeled training data, particularly challenging for 3D medical images where annotation demands substantial domain expertise, time, and resources [1, 2, 3, 4, 5, 6]. Self-Supervised Learning (SSL) has emerged as a transformative solution, offering a powerful mechanism to learn robust, transferable representations from unlabeled data that significantly enhance downstream task performance [7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. This paradigm shift has enabled unprecedented advances in medical image analysis while substantially reducing the annotation burden [2, 4].
In the realm of SSL, Masked Autoencoders (MAE) [17] have demonstrated remarkable success in natural image analysis through their innovative approach to reconstructing distorted views via Vision Transformer architectures [17, 18]. However, extending these methods to 3D medical imaging introduces unique challenges due to the inherent complexity of volumetric data, characterized by high dimensionality and intricate anatomical structures spanning multiple scales. Recent advances have made significant strides in addressing these challenges. MAE3D [19] pioneered the adaptation to 3D medical imaging by reconstructing non-overlapping patches from cropped sub-volumes, establishing a foundation for understanding complex anatomical structures. GL-MAE [20] further advanced this direction by introducing local-global contrastive learning to enhance multi-granular anatomical structure modeling, while SwinUNETR [9] incorporated multi-scale feature learning through hybrid transformers [21]. Although these approaches demonstrated promising results, their reliance on single-scale cropped volumes inherently constrains their ability to capture comprehensive anatomical relationships. Recent efforts have explored alternative strategies to address these limitations: SwinMM [22] leveraged multi-view consistency to enhance volume-wise representations, while Alice [7] utilized pre-trained models [23] to capture rich intra-volume relationships. Despite these advances, three critical challenges persist: the spatial context limitations imposed by volume cropping, which hinder the understanding of complete anatomical structures; the lack of explicit modeling of hierarchical relationships across different anatomical scales; and the substantial computational demands of processing full-resolution 3D volumes. These interconnected challenges underscore the need for a more sophisticated approach that can efficiently capture both fine-grained anatomical details and global contextual information while maintaining practical computational requirements.
@@
Figure 1: Different SSL for 3D medical image analysis. Current Masked Image Modeling methods for 3D medical images primarily (a) rely on pretext tasks e.g., inpainting, at a single level, utilizing hybrid transformers to incorporate all tokens or (b) employ an MAE that reconstructs at a single level using unmasked tokens. In contrast, (c) we observe that 3D medical images inherently exhibit hierarchical properties. Thus, our Mask in Mask (MiM) framework aims to encode multi-level 3D medical image learning across hierarchical visual tokens at various scales through multi-level reconstruction and cross-level alignment (we set the number of level $L$ to 3 in this figure). Additionally, our framework employs a hybrid transformer while only using unmasked tokens.
The intrinsic hierarchical nature of medical images, particularly those with expansive spatial dimensions, necessitates a sophisticated multi-level analytical framework for comprehensive clinical interpretation [24, 9]. As shown in Fig 1, we introduce MiM, a novel hierarchical framework that fundamentally advances MAE-based representation learning for 3D medical images. Our framework systematically addresses the key limitations of existing approaches through three synergistic components. First, to overcome the limited context of cropped volumes, we propose a multi-level volume generation strategy that processes the larger view of 3D volume at multiple scales simultaneously, enabling our model to capture both fine-grained anatomical details and their broader contextual relationships. Second, to explicitly model hierarchical representations, we design a sophisticated multi-level reconstruction mechanism that operates across different anatomical scales. This mechanism preserves critical anatomical details at varying granularities while enforcing consistency through an advanced cross-level alignment strategy, ensuring coherent interpretation between local structures and their global context. Third, to address the computational challenges of processing larger view of the 3D medical images, we incorporate an efficient hybrid backbone design inspired by MCMAE [25]. This architecture significantly reduces computational overhead while maintaining the advantages of transformer-based models, making it practical to analyze high-resolution 3D medical images. These innovations work together synergistically, enabling MiM to effectively model the complete anatomical hierarchy while maintaining computational efficiency.
The principal contributions of this work are threefold:
@@
    We present MiM, a computationally efficient SSL framework that advances MAE through hierarchical design for 3D medical image pre-training. Our approach effectively manages the complexity of 3D medical data while enabling the simultaneous capture of anatomical features across multiple scales, crucial for accurate medical image analysis.
@@
    We introduce a comprehensive methodology for encoding multi-level visual information through two synergistic proxy tasks: multi-level reconstruction and cross-level alignment. This design enables robust local and global representation learning while maintaining anatomical consistency across scales through our novel cross-level alignment mechanism.
@@
    Through extensive experimental validation across twelve diverse datasets, utilizing pre-training sets ranging from 1k to 10k volumes, we demonstrate state-of-the-art performance and establish a clear correlation between pre-training dataset scale and model effectiveness. Our comprehensive evaluation shows significant improvements across various medical imaging tasks, with our multi-level approach consistently outperforming single-scale alternatives.
@@ 2 Related works
Recent advances in SSL for medical image analysis have evolved through different paradigms, with dense prediction tasks like segmentation being crucial yet challenging for 3D medical images. In this section, we first review SSL methods in medical imaging, focusing on the progression from contrastive-based to generative-based approaches and their effectiveness in such dense prediction tasks. This evolution leads to our discussion of Masked Image Modeling, a powerful generative SSL technique showing promise in 3D medical image analysis. However, current Masked Image Modeling approaches struggle to capture the complex hierarchical nature of 3D medical data, motivating our examination of hierarchical SSL designs. Through this review, we demonstrate how existing methods fall short in effectively modeling multi granularity of features in 3D medical images, thereby establishing the context for our proposed hierarchical MIM framework.
SSL for Medical Image Analysis. SSL methods in medical imaging have evolved through two primary paradigms: contrastive-based and generative-based approaches [26, 27, 28]. Contrastive-based methods focus on aligning representations of positive pairs while separating negative pairs in feature space [29]. While foundational works like SimCLR [29], MoCov2 [30], and MoCov3 [31] demonstrated success in medical classification tasks [32, 33], their effectiveness in medical imaging has been further enhanced through domain-specific innovations. For instance,  [34] introduced sub-volume based contrastive learning, while  [35, 23] advanced multi-granular understanding by contrasting features at both local and global scales. However, these approaches’ focus on instance-level alignment has limited their effectiveness in dense prediction tasks like segmentation [32]. Generative-based approaches address these limitations by explicitly modeling spatial structures and preserving local-global consistency [36]. The field has progressed from basic reconstruction tasks using 3D UNet [37, 38] to sophisticated approaches incorporating modern architectures like 3D Swin Transformers with inpainting [9, 34]. This evolution has culminated in Masked Image Modeling, demonstrating robust performance across various 3D medical imaging applications [17, 39, 19, 20, 40].
Medical Image Analysis with Masked Image Modeling. Masked Image Modeling has emerged as a powerful technique for medical image analysis, building upon the success of generative SSL. Initially proposed by [17], this approach demonstrated that high-ratio masking creates an effective self-supervisory task through raw pixel restoration. The field has advanced through innovations in masking strategies [41] and reconstruction targets [18]. While early applications showed promise in 2D medical tasks [42, 43], including disease classification [44] and segmentation [43], the extension to 3D medical imaging introduced new challenges. MAE3D [19] pioneered the adaptation of masked autoencoding to volumetric data, while GL-MAE [20] enhanced anatomical structure understanding by incorporating global-local consistency through contrastive learning. SwinUNETR [9] further advanced the field by recognizing the importance of multi-scale representation learning and incorporating hybrid transformers [21]. However, these methods [19, 20, 9, 22], despite their innovations, process 3D medical images at a single scale with limited receptive fields or rely solely on backbone-level multi-scale processing. As illustrated in Fig. 1, this approach becomes problematic when handling the significantly larger spatial dimensions and varying anatomical structures present in 3D medical images.
Hierarchical SSL. The importance of hierarchical structures in SSL has gained recognition in both natural and medical image domains. In natural image processing, approaches like [18] and [39] have integrated Masked Image Modeling with hierarchical backbones like Swin Transformer [45][46] explored multi-scale feature reconstruction, while [25] advanced masked patch prediction using unmasked contexts. In medical imaging, Adam [47] introduced multi-granular contrastive learning with ResNet [48]. However, these approaches either suffer from the limited dimensionality of 2D images [47, 25] or treat hierarchical levels independently [47], failing to capture the cross-level semantic relationships crucial for 3D medical images.
Distinguished from these approaches, our MiM framework advances the state-of-the-art in 3D medical image analysis through four key innovations: (1) We radically advance beyond MAE3D [19] by designing a powerful multi-level volume generation strategy that enables simultaneous reconstruction of features across both fine and coarse scales, addressing the limitations of existing single-scale methods with limited receptive fields [19, 9, 22, 20] (2) Inspired by  [49, 20], we develop a sophisticated cross-level alignment mechanism specifically for 3D medical volumes that enforces anatomical consistency across hierarchical levels, achieving significant improvements in anatomical feature learning. (3) Our innovative 3D hybrid backbone architecture adapted from 2D natural images [25] achieves superior efficiency in capturing multi-scale features during pre-training while reducing computational demands. (4) MiM demonstrates exceptional scalability by successfully pre-training on massive datasets exceeding 10,000 volumes, surpassing the scale of existing 3D SSL medical imaging methods [38, 32, 19, 20, 9, 8]. The effectiveness of these innovations is validated through extensive experiments across twelve public datasets in organ/tumor segmentation and disease classification tasks, establishing MiM as a transformative advancement in the field.
@@ 3 Methodology
This section provides an overview of our proposed MiM method. Firstly, the overall framework of the MiM method is introduced in Section 3.1. Secondly, the process of multi-level reconstruction is presented in Section 3.2. Then, the cross-level alignment via contrastive learning in our proposed MiM method is described in Section 3.3. Finally, the backbone of the MiM method in the pre-training period is introduced in Section 3.4.
Figure 2: The overall view of our MiM pre-training framework. The level $L$ is set to 3 for better illustration. We first conduct the process of multi-level masked volume generation. The multi-level reconstruction module is responsible for reconstructing the masked volumes at different levels. The cross-level alignment module aligns representations of volumes between volumes from adjacent levels, aiming to enforce anatomical similarity hierarchically.
@@ 3.1 Overall framework
The proposed MiM method’s overall framework is presented in Fig. 2, which comprises multi-level reconstruction modules and cross-level alignment modules. To pre-train the model with MiM, we first generate multi-level volumes from the input 3D medical images. Then, an input volume is cropped into non-overlapping patches, which are divided into unmasked patches and masked patches. The unmasked patches are transformed into a high-dimension feature space using a typical backbone (CNN [48] and transformer [50]). The masked patches are used to generate the next level of masked volumes for multi-level learning. The goal is to restore the masked patches from different levels of masked volumes. In this paper, instead of a single level of masked volume [41, 46], we propose to ease this goal by multi-level of masked volume. We developed a $\mathcal{L}_{R}$ to supervise the final prediction. In addition, we further use a loss function $\mathcal{L}_{C}$ to align the shared semantics between the cross-level volumes, aiming to learn the global semantic content as well as local details. Further details are presented in Section 3.2 and Section 3.3.
@@ 3.2 Multi-level reconstruction
Generation of multi-level masked volumes. Given a volumetric image $x\in\mathbb{R}^{CHWD}$ (e.g., $C=1$ for CT), we aim to generate multi-level volumes $\{x^{l}\in\mathbb{R}^{CH^{l}W^{l}D^{l}},l\in L\}$, which contains $L$ numbers of level, i.e., different granularity information from coarse to fine level. The generation of multi-level volume is illustrated in Fig. 3. We start by cropping a large portion of sub-volume from $x$ as Level-1 volume $x^{1}\in\mathbb{R}^{CH^{1}W^{1}D^{1}}$. The level 1 volume $x^{1}$ is then patchified into $N$ non-overlapping visual tokens using a patch of $\{\frac{H^{1}}{6},\frac{W^{1}}{6},\frac{D^{1}}{6}\}$ as in  [19, 20]. As previous MAE-based methods [20], we apply a high mask ratio $\mu$ (e.g., 60%) to the $N$ non-overlapping tokens (e.g., 216), resulting in unmasked tokens and masked tokens. Each unmasked token is resized at spatial dimension to $\mathbb{R}^{Chwd}$ before further processed by a linear projection as previous MAE-based methods [19]. The masked tokens from Level-1 $x^{1}$ are considered as the next level volume, i.e., Level-2 volume $x^{2}\in\mathbb{R}^{CH^{2}W^{2}D^{2}}$. It’s noted that the Level-2 volume $x^{2}$ is generated from the masked patches of Level-1 volume $x^{1}$, instead of unmasked patches. The content of the target reconstruction at different levels shares regions, meaning that the volumes for Level-1 and Level-2 reconstruct overlapping regions but from different granularities, respectively. This can effectively capture the hierarchical structure of 3D medical images and improve representation learning [46]. Ablation study of the reconstruction targets, i.e., the choices of masked patches or unmasked patches as next level volume, is presented in Section 4.4.1. Since there exist many unmasked tokens, we randomly sample $\gamma$ times without replacement from them for computation efficiency. This generation process of multi-level volume is repeated until the last level of volume $x^{L}$ is generated.
@@
Figure 3: Illustration of the multi-level masked volume generation. Slices drawn from the 3D medical images are shown in the figure only for better illustration. Level-(l+1) volume $x^{l+1}$ is randomly sampled from the masked patches of Level-l volume $x^{l}$ (i.e., patch with red box).
@@
Thus, all unmasked tokens from different levels are fed into the backbone to extract the high-dimension features $z$. Following [17, 19], a lightweight decoder is used to project $z$ along with learnable masked tokens into latent features $q$. For reconstruction, the decoder followed by a simple prediction head is used to reconstruct the masked tokens $y$.
@@
Single-level reconstruction.  Our reconstruction target is the pixel values of the masked tokens for each level volume. With feature extracted from the backbone and the decoder with a prediction head (i.e., a linear layer), following previous MAE methods [19], we reshape the prediction results and reconstruction targets to one-dimensional vector, i.e., $\hat{y}\in\mathbb{R}^{{1\times C}}$ and $y\in\mathbb{R}^{{1\times C}}$, where $C$ is the number of dimensions. Specifically, we empirically set the $C$ to 768, as in  [19, 20].
@@
Then we compute the differences $d$ between the reconstruction target $y$ and the prediction result $\hat{y}$. Specifically, we use MSE distance [19] to measure the difference $d_{m}$ for each masked token $m$ as follows:
@@
where $M$ denotes numbers of masked tokens. To minimize the difference $d_{m}$, we define the reconstruction loss $\mathcal{L}_{\mathcal{R}}^{l}$ for each level of masked volumetric images $l\in L$ as follows:
@@
where $|M|$ represents the number of mask tokens.
Loss function for multi-level reconstruction. To learn multi-granularity details, we apply single-level reconstruction for each level of masked volumetric images. Thus, the multi-level reconstruction loss $\mathcal{L}_{\mathcal{R}}$ can be formulated as follows: